Consider the reaction between 55.0g of copper (II) nitrate tetrahydrate and 55.0 g ammonium phosphate. How many total oxygen atoms participated in the reaction?


1
Expert's answer
2021-04-02T05:15:46-0400

3Cu(NO3)2*4H2O + 2(NH4)3PO4 = Cu3(PO4)2 + NH4NO3 + 12H2O

n((NH4)3PO4) = m/Mr = (55 g)/(149 g/mol) = 0.3691 mol

n(Cu(NO3)2*4H2O) = m/Mr = (55 g)/(260 g/mol) = 0.2115 mol - limiting reactant

Cu(NO3)2*4H2O has 10 oxygen atoms per formula unit and (NH4)3PO4 has 4 oxygen atoms per formula unit, so

n(O) = (0.2115 mol)*(10 + 2/3*4) = 2.679 mol

N(O) = n(O)*NA = (2.679 mol)*(6.02*1023 mol-1) = 1.61*1024 atoms